# On the local thermodynamic relations in relativistic spin hydrodynamics

### v1

@@ Abstract
We demonstrate, by providing two specific examples, that the local differential thermodynamic relations used as educated guesses in relativistic hydrodynamics with spin, do not hold even at global thermodynamic equilibrium, where the expression of the entropy current is totally unambiguous. We show, by using a rigorous quantum statistical method, that for massless free fermions and massive free fermions with rotation and acceleration at global thermodynamic equilibrium, the derivative of the pressure function with respect to the spin potential differs from the spin density and acquires a correction of the same order. Therefore, for an accurate determination of the constitutive relations in relativistic spin hydrodynamics, the traditional method of assuming differential thermodynamic relations is not appropriate.
preprint:
@@ I Introduction
where ^$\widehat{\cal L}^{\lambda\mu\nu}$ is the so-called orbital part and ${\widehat{\cal S}}^{\lambda\mu\nu}$ is the spin tensor. Due to the conservation of $\widehat{\cal J}^{\lambda\mu\nu}$, the divergence of the spin tensor equals the anti-symmetric part of the energy-momentum tensor
@@
As it is well known, the stress-energy and spin tensors are not uniquely defined and one may obtain different pairs of those tensors through a so-called pseudo-gauge transformation [23, 24, 25]
@@
where $\hat{\Pi}$ and $\hat{\Theta}$ are known as super-potentials satisfying, $\hat{\Pi}^{\lambda,\mu\nu}=-\hat{\Pi}^{\lambda,\nu\mu}$ and $\hat{\Theta}^{\mu\nu,\lambda\rho}=-\hat{\Theta}^{\nu\mu,\lambda\rho}=-\hat{\Theta}^{\mu\nu,\rho\lambda}$, respectively. The new pair of stress-energy and spin tensors, however, satisfy the same dynamical equations and do not change the integrated conserved charges. However, both the entropy and entropy production rate are apparently dependent on the choice of the pseudo-gauge [26], thus making the study of the entropy current in spin hydrodynamics an important subject.
Requiring the local entropy production rate to be positive definite, there have been numerous derivations of the constitutive relations of the spin tensor and the stress-energy tensor as a function of the gradients of the thermo-hydrodynamic fields [9, 12, 27, 13, 28, 16, 22, 29, 30]. In those formulations, the entropy current is obtained by assuming the validity of the thermodynamic relations:
@@
However, it was argued in ref. [26] that the differential relation (6) does not generally apply, so that it is not suitable to derive the entropy production rate, whence the constitutive equations of relativistic spin hydrodynamics at the needed level of precision. It was shown that while it is possible to define pressure such that the (5) holds, as well as $\partial p/\partial T|_{\mu,\omega}=s$, it is not guaranteed that:
@@
[…]
@@
which is not guaranteed. The same conclusion applies if $s$ was defined through the (5) and $p$ was known. The only way to check the validity of the above equation, or tantamount (8), is to calculate the thermodynamic functions at the second order in vorticity within a quantum statistical framework.
In this paper, we will do this and we will provide two specific examples of the breaking of the relation (8): for a system of massless free fermions at global thermodynamic equilibrium with rotation and acceleration, for which we know the exact solution; and for a system of massive free fermions at global thermodynamic equilibrium with rotation and acceleration, for which we know the expression of the thermodynamic functions up to second order in vorticity. Since these results are obtained at global thermodynamic equilibrium, where all thermodynamic quantities are unambiguous and no freedom in defining entropy current can be invoked, we conclude that the relation (6) is incorrect, with all due implications.
@@ II Local equilibrium density operator and the entropy current
For a relativistic fluid, by applying the quantum-statistical method, one obtains the local equilibrium density operator, ${\widehat{\rho}}_{LE}$, by maximizing the entropy:
$S=-{\rm Tr}\left({\widehat{\rho}}\log{\widehat{\rho}}\right)\,,$ (9)
@@
where the subscript ‘GE’ stands for global equilibrium. There are other useful relations at global equilibrium involving the acceleration field $a^{\mu}$ and vorticity field $\omega^{\mu}$:
@@
hence, since at global equilibrium $\varpi_{\mu\nu}=\Omega_{\mu\nu}=\omega_{\mu\nu}/T$, we have that the spin potential equals the kinematic vorticity:
$\frac{a^{\mu}}{T}=\frac{1}{T}\omega^{\mu\nu}u_{\nu}\,,\qquad\qquad\frac{\omega^{\mu}}{T}=-\frac{1}{2T}\epsilon^{\mu\nu\rho\sigma}\omega_{\nu\rho}u_{\sigma}\,,$
$\implies\omega_{\mu\nu}=a_{\mu}u_{\nu}-a_{\nu}u_{\mu}+\epsilon_{\mu\nu\rho\sigma}\omega^{\rho}u^{\sigma}\,.$ (19)
@@
$\widehat{\Upsilon}=\frac{1}{T_{0}}\left(\widehat{H}-\mu\widehat{Q}\right)\,,$
@@
In the above equation, the square brackets denote a functional dependence on the thermodynamic fields $T^{\prime},\mu,\omega$. We are now in a position to obtain the entropy current. Starting with Eq. (11) and using Eqs. (20) and (23) we have:
$S$ $=$ $-{\rm Tr}({\widehat{\rho}}_{\rm LE}\log{\widehat{\rho}}_{\rm LE})\,,$ (25)
$=$ $\log Z_{\rm LE}+\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\Big{(}{\rm Tr}({\widehat{\rho}_{\rm LE}}{\widehat{T}}^{\mu\nu})\beta_{\nu}-\zeta{\rm Tr}({\widehat{\rho}_{\rm LE}}{\widehat{j}}^{\mu})$
$-\frac{1}{2}\Omega_{\lambda\nu}{\rm Tr}({\widehat{\rho}_{\rm LE}}{\widehat{\cal S}}^{\mu\lambda\nu})\Big{)}\,,$
$=$ $\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\left(\phi^{\mu}+T^{\mu\nu}_{\rm LE}\beta_{\nu}-\zeta j^{\mu}_{\rm LE}-\frac{1}{2}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}_{\rm LE}\right)$
@@ III Local thermodynamic relations for a free Dirac field at global equilibrium
which is the first thermodynamic relation in eq. (5). Following this definition and using the eq. (28) we have:
@@
-- $\frac{1}{2}\omega_{\lambda\nu}u_{\mu}\mathcal{S}^{\mu\lambda\nu}[T^{\prime}]\Big{)}$
@@
$\frac{\partial p}{\partial T}\Big{|}_{\mu,\omega}=s\,,$
@@ III.1 Massless Dirac fermions
The exact mean values of the currents normalized so as to vanish in the Minkowski vacuum $|0_{M}\rangle$, at global thermodynamic equilibrium with the density operator (17) have been derived in ref. [34, 35] for the massless Dirac field. In order not to confuse them with the mean values normalized so as to vanish in the lowest lying state of the operator $\widehat{\Upsilon}$, that is, in the limit $T\to 0$, we will denote them by a tilde, hence:
@@
Applying equation (32), we get the following simple forms for the vector and axial currents to be used in the entropy current calculation:
@@
In refs. [34, 35] only the exact expression of the symmetric Belinfante stress-energy tensor was derived. However, since we need to tackle thermodynamic relations with a spin tensor, we have to select a pseudo-gauge where the spin tensor does not vanish. For the free Dirac field, the simplest pseudo-gauge with a non-trivial spin tensor is the canonical one, which is related to the Belinfante pseudo-gauge by the following equation:
@@
$u_{\mu}u_{\nu}T^{\mu\nu}_{\rm Can}=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm B}-\frac{1}{2}u_{\mu}u_{\nu}\partial_{\lambda}\mathcal{S}^{\lambda\mu\nu}_{\rm Can}$ (37)
$=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm B}=\rho=\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}-\frac{a^{2}T^{2}}{24}-\frac{\omega^{2}T^{2}}{8},$
@@ III.2 Massive Dirac fermions
A similar calculation can be carried out for massive fermions at second order of the expansion in thermal vorticity at global thermodynamic equilibrium. The general expression, up to that order, of the stress-energy tensor, the vector and axial currents was obtained in ref. [39]:
$T^{\mu\nu}_{\rm B}$ $=$ $\left(\rho-\frac{a^{2}U_{\alpha}}{T^{2}}-\frac{\omega^{2}U_{w}}{T^{2}}\right)u^{\mu}u^{\nu}+\frac{A}{T^{2}}a^{\mu}a^{\nu}$ (46)
-- $\left(p-\frac{a^{2}D_{\alpha}}{T^{2}}-\frac{\omega^{2}D_{w}}{T^{2}}\right)\Delta^{\mu\nu}+\frac{W}{T^{2}}\omega^{\mu}\omega^{\nu}$
@@
$j^{\mu}$ $=$ $n\,u^{\mu}+\left(\frac{a^{2}N_{\alpha}}{T^{2}}+\frac{\omega^{2}N_{w}}{T^{2}}\right)u^{\mu}+G_{V}l^{\mu}\,,$ (47)
$j^{\mu}_{A}$ $=$ $\frac{1}{T}\omega^{\mu}W^{A}\,.$ (48)
The expression of the coefficients $\rho$, $p$, $U_{\alpha}$, $U_{w}$, $A$, $D_{\alpha}$, $D_{w}$, $W$, $G$, $n$, $N_{\alpha}$, $N_{w}$, $G_{V}$, and $W_{A}$ is reported in Appendix B in the equation (62). By using the equations (35) and (36) we find:
$u_{\mu}u_{\nu}T^{\mu\nu}_{\rm Can}$ $=$ $\rho-\frac{a^{2}U_{\alpha}}{T^{2}}-\frac{\omega^{2}U_{w}}{T^{2}}\,,$
$\mu u_{\mu}j^{\mu}$ $=$ $\mu\,n+\frac{\mu}{T^{2}}\left(a^{2}N_{\alpha}+\omega^{2}N_{w}\right)\,,$
$u_{\mu}\omega_{\lambda\nu}\mathcal{S}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\frac{1}{T}\omega^{2}W^{A}\,.$ (49)
@@
$\rho$ $=$ $\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}+\frac{\mu^{4}}{4\pi^{2}},\quad U_{\alpha}=\left(\frac{T^{4}}{24}+\frac{\mu^{2}T^{2}}{8\pi^{2}}\right),$
@@
$n$ $=$ $\frac{T^{2}\mu}{3}\left(1+\frac{\mu^{2}}{\pi^{2}T^{2}}\right)\,,\quad N_{\alpha}=N_{w}=\frac{T^{2}\mu}{4\pi^{2}}\,,$
thus, eqs. (49) become:
$u_{\nu}u_{\mu}T^{\mu\nu}_{\rm Can}$ $=$ $\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}-\frac{a^{2}}{24}T^{2}-\frac{\omega^{2}}{8}T^{2},$
$\mu u_{\mu}j^{\mu}$ $=$ $\frac{T^{2}\mu^{2}}{3},$
@@
After subtracting the limits for $T\to 0$, we can obtain the pressure function of eq. (III), which fully matches eq. (42).
@@
$\rho$ $=$ $\frac{m^{4}}{\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
@@
$n$ $=$ $\frac{m^{3}}{\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
@@
$u_{\nu}u_{\mu}T^{\mu\nu}_{\rm Can}$ $=$ $\frac{m^{4}{\rm e}^{-(x-\zeta)}}{\sqrt{2}(\pi x)^{3/2}}\left(1-\frac{a^{2}x}{24\,T^{2}}-\frac{\omega^{2}}{8\,T^{2}}\right),$
$\mu u_{\mu}j^{\mu}$ $=$ $\frac{\mu\,m^{3}{\rm e}^{-(x-\zeta)}}{\sqrt{2}(\pi x)^{3/2}}\left(1+\frac{a^{2}x}{24\,T^{2}}+\frac{\omega^{2}}{8\,T^{2}}\right),$
$u_{\mu}\omega_{\lambda\nu}\mathcal{S}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\frac{m^{3}\,\omega^{2}}{2\sqrt{2}\,T(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}.$ (54)
Therefore, in the non-relativistic limit, the pressure $p$ defined in eq. (III) reads:
@@
IV Conclusions
In conclusion, we have shown, by providing two specific counter-examples, namely a system of massless free fermions and massive free fermions with rotation and acceleration at global thermodynamic equilibrium, that the differential thermodynamic relations (6) often used to obtain the constitutive equations in relativistic spin hydrodynamics do not hold. Notably, the derivative of the pressure function with respect to spin potential (which reduces to vorticity at global equilibrium) features additional contributions of the same order of magnitude as the spin density. The relations (5) and (6) therefore can hardly be used to make a precise derivation of the constitutive relations in relativistic spin hydrodynamics; their use may easily lead to missing some dissipative terms for the spin tensor. Indeed, the quantum-statistical method proposed in refs. [26, 32] to determine the entropy production rate and constitutive equations in relativistic spin hydrodynamics led to the identification of several new dissipative coefficients for the spin tensor.

### v3

@@ Abstract
We demonstrate, by providing two specific examples, that the local differential thermodynamic relations used as educated guesses in relativistic hydrodynamics with spin, do not hold even at global thermodynamic equilibrium. We show, by using a rigorous quantum statistical method, that for massless free fermions and massive free fermions with rotation and acceleration at global thermodynamic equilibrium, the derivative of the pressure function with respect to the spin potential differs from the spin density and acquires a correction of the same order. Such correction cannot be eliminated by any redefinition of the entropy current, a so-called entropy-gauge transformation. Therefore, for an accurate determination of the constitutive relations in relativistic spin hydrodynamics, the traditional method of assuming differential thermodynamic relations is not appropriate.
@@ I Introduction
where $\widehat{\cal L}^{\lambda\mu\nu}$ is the so-called orbital part and ${\widehat{\cal S}}^{\lambda\mu\nu}$ is the spin tensor. Due to the conservation of $\widehat{\cal J}^{\lambda\mu\nu}$, the divergence of the spin tensor equals the anti-symmetric part of the energy-momentum tensor
@@
As it is well known, the energy-momentum and spin tensors are not uniquely defined and one may obtain different pairs of those tensors through a so-called pseudo-gauge transformation [23, 24, 25]
@@
where $\hat{\Pi}$ and $\hat{\Theta}$ are known as super-potentials satisfying, $\hat{\Pi}^{\lambda,\mu\nu}=-\hat{\Pi}^{\lambda,\nu\mu}$ and $\hat{\Theta}^{\mu\nu,\lambda\rho}=-\hat{\Theta}^{\nu\mu,\lambda\rho}=-\hat{\Theta}^{\mu\nu,\rho\lambda}$, respectively. The new pair of energy-momentum and spin tensors, however, satisfy the same dynamical equations and do not change the integrated conserved charges. However, both the entropy and entropy production rate are apparently dependent on the choice of the pseudo-gauge [26], thus making the study of the entropy current in spin hydrodynamics an important subject.
Requiring the local entropy production rate to be positive definite, there have been numerous derivations of the constitutive relations of the spin tensor and the energy-momentum tensor as a function of the gradients of the thermo-hydrodynamic fields [9, 12, 27, 13, 28, 16, 22]. In those formulations, the entropy current is obtained by assuming the validity of the thermodynamic relations:
@@
However, for a given choice of the pseudo-gauge, it was argued in ref. [26] that the differential relation (6) does not generally apply, so that it is not suitable to derive the entropy production rate, whence the constitutive equations of relativistic spin hydrodynamics at the needed level of precision. The inadequacy of the equation (6) was also the subject of a recent study [29, 30] where the authors proposed to replace it with a differential relation for vector currents, which we will not delve into in this work.
In ref. [26] it was shown that while it is possible to define pressure such that the (5) holds, as well as $\partial p/\partial T|_{\mu,\omega}=s$, it is not guaranteed that:
@@
The argument is based on a first-principle derivation of the entropy current within a quantum-relativistic statistical mechanics framework proposed in ref. [31] and this method was further applied to derive the entropy production rate and the constitutive equations in relativistic spin hydrodynamics [32]. It can be figured out that a simple addition of the term $\frac{1}{2}\omega_{\mu\nu}S^{\mu\nu}$ in the usual thermodynamic relation (5) does not necessarily imply the validity of the eq. (8) with a simple argument, which applies at global equilibrium, where the reduced spin potential is equal to the thermal vorticity tensor $\varpi$, as we will see. The spin density $S^{\mu\nu}$ is, at the leading order, first order in vorticity, so in order to test the (6) one needs to know all thermodynamic functions $s$, $\rho$, $p$, $n$, and $S^{\mu\nu}$ at least up to second order in vorticity. Even if we define one of them, e.g. $p$ or $s$, by using the (5), for the (8) to hold we would need that the independently calculated thermodynamic functions fulfill a special constraint. For instance, in case of using (5) to define $p$:
@@
which is not guaranteed. The same conclusion applies if $s$ was defined through the (5) and $p$ was known.
In this paper, we will check the validity of the equation (8) by calculating the thermodynamic functions at the second order in vorticity within a quantum statistical framework and obtain two specific instances of the breaking of the relation (8): for massless free fermions at global thermodynamic equilibrium with rotation and acceleration in the canonical pseudo-gauge, for which the exact solution is known; and for a system of massive free fermions at global thermodynamic equilibrium with rotation and acceleration, again, in the canonical pseudo-gauge, for which we know the expression of the thermodynamic functions up to second order in vorticity. We will further show that, in the examined cases, no redefinition of the entropy current can restore the differential relations (6) and so we will conclude that at least one of them is incorrect, with all due implications.
@@ II Local equilibrium density operator and the entropy current
For a relativistic fluid, by applying the quantum-statistical method, one obtains the local equilibrium density operator, ${\widehat{\rho}}_{\rm LE}$, by maximizing the entropy:
[…]
@@
where the subscript ‘GE’ stands for global equilibrium and $^{\widehat{P}}$, $^{\widehat{J}}$, and $^{\widehat{Q}}$ are the four-momentum operator, boost-angular momentum operator, and conserved charge, respectively. There are other useful relations at global equilibrium involving the acceleration field $a^{\mu}$ and vorticity field $\omega^{\mu}$:
@@
hence, since at global equilibrium $\varpi_{\mu\nu}=\Omega_{\mu\nu}=\omega_{\mu\nu}/T$, we have a very useful formula for the spin potential at global equilibrium:
$\omega_{\mu\nu}=a_{\mu}u_{\nu}-a_{\nu}u_{\mu}+\epsilon_{\mu\nu\rho\sigma}\omega^{\rho}u^{\sigma}\,.$ (19)
@@
[…]
@@
In the above equation, the square brackets denote a functional dependence on the thermodynamic fields $T^{\prime}$, $\mu$, and $\omega$. We are now in a position to obtain the entropy current. Starting with Eq. (11) and using Eqs. (20) and (23) we have:
$S$ $=$ $-{\rm Tr}\left({\widehat{\rho}}_{\rm LE}\log{\widehat{\rho}}_{\rm LE}\right)\,$, (25)
$=$ $\log Z_{\rm LE}+\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\Big({\rm Tr}\left({\widehat{\rho}_{\rm LE}}{\widehat{T}}^{\mu\nu}\right)\beta_{\nu}-\zeta{\rm Tr}\left({\widehat{\rho}_{\rm LE}}{\widehat{j}}^{\mu}\right)$
$-\frac{1}{2}\Omega_{\lambda\nu}{\rm Tr}\left({\widehat{\rho}_{\rm LE}}{\widehat{\cal S}}^{\mu\lambda\nu}\right)\Big)\,$,
$=$ $\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\left(\phi^{\mu}+T^{\mu\nu}_{\rm LE}\beta_{\nu}-\zeta j^{\mu}_{\rm LE}-\frac{1}{2}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}_{\rm LE}\right)\,$,
@@ III Local thermodynamic relations for a free Dirac field at global equilibrium
which is the first thermodynamic relation in eq. (5). It should be emphasized that the hereby defined pressure does not coincide in general with the average diagonal stress in the energy-momentum tensor. Particularly, it does not need to fulfill the equation $p=\rho/3$ for a massless Dirac field.
Following this definition and using the eq. (28) we have:
@@
- $\frac{1}{2}\omega_{\lambda\nu}u_{\mu}\mathcal{S}^{\mu\lambda\nu}[T^{\prime}]\Big)\,$,
@@
$\frac{\partial p}{\partial T}\Big|_{\mu,\omega}=s\,$, (32)
@@ III.1 Massless Dirac fermions
The exact mean values of the currents normalized so as to vanish in the Minkowski vacuum $|0_{M}\rangle$, at global thermodynamic equilibrium with the density operator (17) have been derived in refs. [34, 35] for the massless Dirac field. In order not to confuse them with the mean values normalized so as to vanish in the lowest lying state of the operator $\widehat{\Upsilon}$, that is, in the limit $T\to 0$, we will denote them by a tilde, hence:
@@
Applying equation (33) to subtract the zero temperature limit, we get the following simple forms for the vector and axial currents to be used in the entropy current calculation:
@@
In refs. [34, 35] only the exact expression of the symmetric Belinfante energy-momentum tensor was derived. However, since we need to tackle thermodynamic relations with a spin tensor, we have to select a pseudo-gauge where the spin tensor does not vanish. For the free Dirac field, the simplest pseudo-gauge with a non-trivial spin tensor is the canonical one, which is related to the Belinfante pseudo-gauge by the following equation:
@@
$\rho=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm Can}=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm B}-\frac{1}{2}u_{\mu}u_{\nu}\partial_{\lambda}\mathcal{S}^{\lambda\mu\nu}_{\rm Can}$ (38)
$=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm B}=\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}-\frac{a^{2}T^{2}}{24}-\frac{\omega^{2}T^{2}}{8},$
@@ III.2 Massive Dirac fermions
A similar calculation can be carried out for massive fermions at second order of the expansion in thermal vorticity at global thermodynamic equilibrium. The general expression, up to that order, of the Belinfante energy-momentum tensor, the vector and axial currents normalized with Minkowski vacuum subtraction was obtained in ref. [39]:
$\tilde{T}^{\mu\nu}_{\rm B}$ $=$ $\left(\bar{\rho}-\frac{a^{2}U_{\alpha}}{T^{2}}-\frac{\omega^{2}U_{w}}{T^{2}}\right)u^{\mu}u^{\nu}+\frac{A}{T^{2}}a^{\mu}a^{\nu}$ (47)
- $\left(\bar{p}-\frac{a^{2}D_{\alpha}}{T^{2}}-\frac{\omega^{2}D_{w}}{T^{2}}\right)\Delta^{\mu\nu}+\frac{W}{T^{2}}\omega^{\mu}\omega^{\nu}$
@@
$\tilde{j}^{\mu}$ $=$ $\bar{n}\,u^{\mu}+\left(\frac{a^{2}N_{\alpha}}{T^{2}}+\frac{\omega^{2}N_{w}}{T^{2}}\right)u^{\mu}+G_{V}l^{\mu}\,$, (48)
$\tilde{j}^{\mu}_{A}$ $=$ $\frac{1}{T}\omega^{\mu}W^{A}\,.$ (49)
The expression of the coefficients $\bar{\rho}$, $\bar{p}$, $U_{\alpha}$, $U_{w}$, $A$, $D_{\alpha}$, $D_{w}$, $W$, $G$, $\bar{n}$, $N_{\alpha}$, $N_{w}$, $G_{V}$, and $W_{A}$ is reported in Appendix B in the equation (69). By using the equations (36) and (37) we find:
$u_{\mu}u_{\nu}\tilde{T}^{\mu\nu}_{\rm Can}$ $=$ $\bar{\rho}-\frac{a^{2}U_{\alpha}}{T^{2}}-\frac{\omega^{2}U_{w}}{T^{2}}\,$,
$u_{\mu}\tilde{j}^{\mu}$ $=$ $\bar{n}+\frac{1}{T^{2}}\left(a^{2}N_{\alpha}+\omega^{2}N_{w}\right)\,$,
$u_{\mu}\omega_{\lambda\nu}\tilde{\mathcal{S}}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\frac{1}{T}\omega^{2}W^{A}\,.$ (50)
@@
$\bar{\rho}$ $=$ $\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}+\frac{\mu^{4}}{4\pi^{2}},\quad U_{\alpha}=\left(\frac{T^{4}}{24}+\frac{\mu^{2}T^{2}}{8\pi^{2}}\right),$
@@
$\bar{n}$ $=$ $\frac{T^{2}\mu}{3}\left(1+\frac{\mu^{2}}{\pi^{2}T^{2}}\right)\,,\quad N_{\alpha}=N_{w}=\frac{T^{2}\mu}{4\pi^{2}}\,.$
After subtraction of the zero temperature limit as prescribed, eqs. (50) give rise to:
$u_{\nu}u_{\mu}T^{\mu\nu}_{\rm Can}$ $=$ $\rho=\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}-\frac{a^{2}}{24}T^{2}-\frac{\omega^{2}}{8}T^{2},$
$u_{\mu}j^{\mu}$ $=$ $n=\frac{T^{2}\mu}{3},$
@@
The ensuing pressure function defined in eq. (III) fully matches the one in eq. (43).
@@
$\bar{\rho}$ $=$ $\frac{m^{4}}{\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,$,
@@
$\bar{n}$ $=$ $\frac{m^{3}}{\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,$,
@@
$u_{\nu}u_{\mu}\tilde{T}^{\mu\nu}_{\rm Can}$ $=$ $\frac{m^{4}{\rm e}^{-(x-\zeta)}}{\sqrt{2}(\pi x)^{3/2}}\left(1-\frac{a^{2}x}{24\,T^{2}}-\frac{\omega^{2}}{8\,T^{2}}\right),$
$u_{\mu}\tilde{j}^{\mu}$ $=$ $\frac{m^{3}{\rm e}^{-(x-\zeta)}}{\sqrt{2}(\pi x)^{3/2}}\left(1+\frac{a^{2}x}{24\,T^{2}}+\frac{\omega^{2}}{8\,T^{2}}\right),$
$u_{\mu}\omega_{\lambda\nu}\tilde{\mathcal{S}}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\frac{m^{3}\,\omega^{2}}{2\sqrt{2}\,T(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}.$ (55)
The zero temperature limit vanishes for all quantities in (III.2), so that the thermodynamic functions $\rho$, $n$ and $\omega_{\lambda\nu}S^{\lambda\nu}$ coincide with the right-hand side of the above equations. Therefore, in the non-relativistic limit, the pressure $p$ defined in eq. (III) reads:
@@
IV Entropy-gauge transformations
There is a certain freedom in defining the entropy current $s^{\mu}$. All in all, its main purpose is to yield the total entropy $S$ by integrating it over a space-like 3D hypersurface:
$S=\int_{\Sigma}{\rm d}\Sigma_{\mu}\;s^{\mu}\;.$
At global thermodynamic equilibrium, the total entropy $S$ is independent of the hypersurface $\Sigma$, which requires that the entropy current has zero divergence, $\partial_{\mu}s^{\mu}=0$. That being said, the entropy current can be modified by adding the divergence of an anti-symmetric field $A^{\lambda\mu}$:
$s^{\prime\mu}=s^{\mu}+\partial_{\lambda}A^{\lambda\mu}\,$, (61)
which leaves its divergence invariant as well as the total entropy provided that suitable boundary conditions for the field $A$ are set. The transformation (61) of the entropy current has been named entropy-gauge transformation in ref. [26]. Indeed, one may wonder whether through an entropy-gauge transformation the traditional relations (5) and (6) can be restored. To begin with, the thermodynamic potential current $\phi^{\mu}$ can be modified likewise so as to maintain the relation (23) as well as the (27); in symbols:
$\phi^{\prime\mu}=\phi^{\mu}+\partial_{\lambda}A^{\lambda\mu}\,$, (62)
With the (61) and (62) the (27) is invariant and can be rewritten as:
$s^{\prime\mu}=\phi^{\prime\mu}+T^{\mu\nu}\beta_{\nu}-\zeta j^{\mu}-\frac{1}{2}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}\,.$ (63)
To keep the local thermodynamic relation in the form of (5) one must redefine the pressure $p^{\prime}=Tu_{\mu}\phi^{\prime\mu}$. The question is whether with such a new definition it is possible to implement the (6), notably:
$\frac{\partial p^{\prime}}{\partial T}\Big|_{\mu,\omega}=s^{\prime}\,,\qquad\frac{\partial p^{\prime}}{\partial\omega_{\lambda\nu}}\Big|_{T,\mu}=S^{\lambda\nu}\,.$
Taking into account the equations (32) and (61), the first equation on the left requires [26]:
$\frac{\partial}{\partial T}\left(u_{\mu}\partial_{\lambda}A^{\lambda\mu}\right)\Big|_{\mu,\omega}\equiv\frac{\partial\Gamma}{\partial T}\Big|_{\mu,\omega}=0\,$, (64)
while the second:
$\frac{\partial\Gamma}{\partial\omega_{\lambda\nu}}\Big|_{T,\mu}=S^{\lambda\nu}-\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big|_{T,\mu}\,.$ (65)
The next question is whether such a function $\Gamma$ exists. By taking the derivative of the (65) with respect to $T$ and of the (64) with respect to $\omega_{\lambda\nu}$, we derive a necessary condition for its existence, namely:
$\frac{\partial}{\partial T}\left(S^{\lambda\nu}-\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big|_{T,\mu}\right)\Big|_{\mu,\omega}=0\,$, (66)
which is equivalent to the second-order mixed derivative condition implied by the (6):
$\frac{\partial S^{\lambda\nu}}{\partial T}\Big|_{\mu,\omega}=\frac{\partial s}{\partial\omega_{\lambda\nu}}\Big|_{T,\mu}\,.$
In the massless fermion case, according to the equation (III.1) we have:
$S^{\lambda\nu}-\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big|_{T,\mu}$ $=\frac{T^{2}}{12}(a^{\lambda}u^{\nu}-a^{\nu}u^{\lambda})\,$,
$=\frac{T^{2}}{12}(\omega^{\lambda\mu}u_{\mu}u^{\nu}-\omega^{\nu\mu}u_{\mu}u^{\lambda})\,$,
which does not manifestly fulfill equation (66). The same conclusion is achieved for the massive fermion case, by using the equations (III.2) and (59). Therefore, we conclude that a redefinition of the entropy current cannot restore the differential thermodynamic relations (6).
V Conclusions
In conclusion, we have shown, by providing two specific counter-examples, namely a system of massless free fermions and massive free fermions with rotation and acceleration at global thermodynamic equilibrium in the canonical pseudo-gauge, that the differential thermodynamic relations (6) often used to obtain the constitutive equations in relativistic spin hydrodynamics do not hold. Notably, the derivative of the pressure function with respect to spin potential (which reduces to vorticity at global equilibrium) features additional contributions of the same order of magnitude as the spin density. The differential thermodynamic relations cannot be restored by a suitable redefinition of the entropy current (an entropy-gauge transformation).
The relations (5) and (6) therefore can hardly be used to make a precise derivation of the constitutive relations in relativistic spin hydrodynamics; their use may easily lead to missing some dissipative terms for the spin tensor. Indeed, the quantum-statistical method proposed in refs. [26, 32] to determine the entropy production rate and constitutive equations in relativistic spin hydrodynamics led to the identification of several new dissipative coefficients for the spin tensor.